A company is analyzing two mutually exclusive projects, S and L, with the following cash flows:

0 1 2 3 4
Project S -$1,000 $884.04 $250 $10 $10
Project L -$1,000 $10 $260 $420 $797.68

The company's WACC is 9.0%. What is the IRR of the better project? (Hint: The better project may or may not be the one with the higher IRR.) Round your answer to two decimal places.
Answer #1

S:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=884.04/1.09+250/1.09^2+10/1.09^3+10/1.09^4

=1036.27

NPV=Present value of inflows-Present value of outflows

=1036.27-1000

=$36.27(Approx)

L:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=10/1.09+260/1.09^2+420/1.09^3+797.68/1.09^4

=1117.42

NPV=Present value of inflows-Present value of outflows

=1117.42-1000

=$117.42(Approx)

Hence L must be selected having higher NPV.

Let irr be x%
At irr,present value of inflows=present value of outflows.

1000=10/1.0x+260/1.0x^2+420/1.0x^3+797.68/1.0x^4

Hence x=irr=12.75%(Approx).